OpenStudy (anonymous):

x^2 - 2k(x+1) +k^2 = 6 has real roots. Find the roots in terms of k. Ans given: sqrt(2k+6) and - sqrt(2k+6)

OpenStudy (anonymous):

is this a repeat?

OpenStudy (anonymous):

Since the equation has real roots, then : \[\Delta\geq0\] The standard form of the equation : \[x^2-2kx+k^2-2k-6=0\] Then : \[\Delta=4k^2-4\times(k^2-2k-6)=8k+24=8(k+3)\] So the roots \[x_1=\frac{2k+2\sqrt{2(k+3)}}{2}=k+\sqrt{2(k+3)} \\x_2=k-\sqrt{2(k+3)} \]

OpenStudy (anonymous):

yeah i was totally wrong

OpenStudy (anonymous):

@Noura11 I dont understand how you got to Δ=4k^2−4×(k^2−2k−6) sorry :/

OpenStudy (anonymous):

OK ! The standard form of the equation : \[x^2−2kx+k^2−2k−6=0\] Then : \[a=1~~ b=-2k~~~ c=k^2-2k+6\] And we know : \[\Delta= b^2-4\times a\times c\]